# Question-Analysis Prompting Improves LLM Performance in Reasoning Tasks

### v1

@@ Abstract
Although LLMs have the potential to transform many fields, they still underperform humans in reasoning tasks. Existing methods induce the model to produce step-by-step calculations, but this research explores the question: Does making the LLM analyze the question improve its performance? We propose a novel prompting strategy called Question Analysis Prompting (QAP), in which the model is prompted to explain the question in n words before solving. The value of n influences the length of response generated by the model. QAP is evaluated on GPT 3.5 Turbo and GPT 4 Turbo on arithmetic datasets GSM8K, AQuA, and SAT and commonsense dataset StrategyQA. QAP is compared with other state-of- the-art prompts including Chain-of-Thought (CoT), Plan and Solve Prompting (PS+) and Take A Deep Breath (TADB). QAP outperforms all state-of-the-art prompts on AQuA and SAT datasets on both GPT3.5 and GPT4. QAP consistently ranks among the top-2 prompts on 75% of the tests. A key factor of QAP performance can be attributed to response length, where detailed responses are beneficial when answering harder questions, but can negatively affect easy questions.
@@ 1 Introduction
Large language models (LLMs) have recently shown rapid improvement across a host of standard natural language processing (NLP) tasks, including arithmetic, commonsense and symbolic reasoning. (Brown et al., 2020) Although these models show improved ability to understand and generate text (OpenAI, 2023), their performance can still be further improved. One solution is to encourage the model to think step-by-step. Using chain-of-thought prompting (Wei et al., 2022), LLMs are given Q&A exemplars which are designed to elicit a structured step-by-step response from the model. Many newly developed strategies meant to improve LLM performance have been focused on sophisticating the model’s step-by-step calculation (Gu et al., 2023). Despite SoTA prompts’ remarkable success across various tasks, their accuracies can still be further improved. In this work, we explore ways to improve the model reasoning not only in the answer steps, but also how the model interprets the question itself. By making the model to explicitly interpret the question, we maximize its understanding of the question and minimize missed key information. This paper introduces Question-Analysis Prompting (QAP), a simple zero-shot prompting strategy that induces the model to first explain the question before solving. This method is adaptable to various problem difficulties and shows promising results in math and commonsense reasoning across different model sizes.
@@ 2 Prompt Design
The key principle behind QAP is that the model should reiterate the problem before solving. Another principle is that we should be able to control how much the model explains so that we can adapt the prompt to different model sizes and problem complexities. The specific prompt used is as follows:
@@ 3 Prompt Impact
In Figure 1, we highlight the structure of a standard QAP output. First, the model breaks down the question in its own words and provides in-depth analysis on each event. We notice a direct relationship between the explanation and the answer steps. Each calculation is previously mentioned in the explanation portion, and this proves that the explanation has allowed the model to plan its approach even before solving. As a result, there is a significant increase in step-by-step calculation and a decreased chance of missed steps.
@@ 4.1 Benchmarks
We evaluate the effectiveness of QAP on three arithmetic reasoning datasets. These include grade-school math questions from GSM8K (Cobbe et al., 2021), algebraic word problems from AQuA (Ling et al., 2017), and SAT math problems from AGIEval (Zhong et al., 2023). For commonsense reasoning, we evaluate on open-domain questions that require implicit reasoning, from StrategyQA (Geva et al., 2021). We evaluate on the test sets of all benchmarks, as some proprietary models are partially trained on the training set of such tasks. (OpenAI, 2023)
@@ 4.2 Models
We specifically choose our models to observe the prompts’ impacts across differences in model size. The smaller model is GPT3.5 Turbo with version gpt-3.5-turbo-0613. Our larger model is GPT4 Turbo with version gpt-4-1106-preview (OpenAI, 2023). For both of the models we used the OpenAI API  [^1] https://platform.openai.com/docs/api-reference/chat for running our experiments.
@@ 4.3 Prompts
For all datasets and models, we experiment with different variations of QAP. We utilize QAP25, QAP50, QAP100, QAP150, and QAP200. We compare the performance of QAP with the baseline (no prompt). Additionally we compare QAP with two different zero-shot prompts, TADB - "Take a deep breath and work on this problem step-by-step" (Yang et al., 2023) and PS+ (Plan and Solve Plus) (Wang et al., 2023). Finally we also compare QAP with 8-shot chain-of-thought prompting.
@@ 4.4 Results
The results for GPT-3.5 and GPT-4 Turbo are shown in Table  Table 1 and Table  Table 2 respectively. General word counts are shown in Figure 7.
Arithmetic Reasoning: On GPT 3.5 Turbo, a variant of QAP is the top performer in 2 out of 3 arithmetic tasks. QAP shows significant gains on AQuA and SAT. With GPT-4 Turbo, QAP performs the best in the same 2 out of 3 arithmetic tasks. This suggests that QAP may be more beneficial on questions involving algebraic and higher-level problem solving; additionally, GPT-4 is trained at least in part on GSM8K (OpenAI, 2023) and thus its performance may be less sensitive to prompting changes.
@@
Table 1: Results for GPT-3.5 Turbo (highest scores bolded)
@@
Table 2: Results for GPT 4 Turbo. (highest scores bolded)
Commonsense Reasoning:. On StrategyQA, QAP consistently performs second-best when compared to other prompts. On both models, QAP25 is the highest QAP performer. This suggests that fewer-word explanations benefit commonsense reasoning. This is because too much explanation can cause the model to confuse a simple answer 6 While there is a decline in performance as $n$ increases on the 3.5 model, the larger GPT-4 Turbo model yields similar performances across all QAP variants.
@@ 5 Analysis
Question Difficulties Based On Baseline Performance: Within a given dataset, the difficulty of the individual question may vary. We propose a method to measure question difficulty based on performance with the baseline prompt. If the model can answer the problem correctly with the baseline prompt, then we consider the question to be easy; otherwise the question is hard. We analyze the performance of different prompts across “easy” and “hard” questions. Table 3 and Table 4 shows that QAP consistently outperforms other prompts in the “hard” category.
Impact Of Word Counts On Question Difficulties: QAP generates higher word counts for both “easy" and “hard" questions ( Table 5 and  Table 6 ), despite performing lower on “easy” questions. Although more step-by-step thought processes are encouraged to avoid mistakes during reasoning, this suggests that over-explanation can negatively impact the mode (also shown in Figure  Figure 5). Thus, the most suitable word count to solve a problem will vary from task to task; longer explanations are best suited to more complicated questions for which baseline prompting fails.
Downsides Of Smaller QAPs: Despite high performance on StrategyQA, QAP25 performs poorly on arithmetic datasets (mostly SAT and AQuA) using GPT-3.5 Turbo. Due to a small value of n, the model outputs are unfinished responses (i.e. the model stops midway through its reasoning steps) (shown in Figure 8) On SAT math, 51% of responses were incomplete for QAP25. On AQuA, 19% of responses were incomplete for QAP25.
Figure 2: We consider difficulty of the problem based on baseline’s results. E.g., an incorrect answer is “hard” and a correct answer is “easy”. Left chart shows accuracy within each difficulty. Right chart shows mean (average) word count for within each difficulty. All results for each prompt are shown in Table: 6 and Table:4
@@ 6 Additional Studies
Placement of the prompt: In this evaluation, we studied the impact of prompt placement on performance using GSM8K dataset. Two options for prompt placement were considered, Q_Begin - adding the prompt before the question and Q_End - adding the prompt after the question. Both placements provided similar results on GPT-3.5 and GPT-4. Results shown in the rest of the paper are based on Q_End.
Two-stage QAP: In this approach, we performed the prompting in two stages. In the first stage the model is prompted with, “Explain this problem to me in at least 50 words WITHOUT SOLVING.” In the second stage, the model is prompted again with the question and the explanation from the first stage. On GSM8K and AQuA, the model not only explained the problem, but also outlined steps needed to solve it. However, the accuracy was almost 50% worse than single stage prompting.
@@ 7 Related Work
Chain of Thought: Chain-of-thought reasoning is a notable few-shot (zero-shot also exists (Yang et al., 2023) example in which the model is shown how to express its reasoning steps (Wei et al., 2022). This approach was highly effective as the model would replicate these exemplars, and their accuracies improved drastically. CoT encouraged the model to think step-by-step, and this concept would be repeating theme among other zero-shot counterparts.
@@
Plan and Solve Prompting Plus: Another zero-shot prompt is Plan-and-Solve Prompting (Wang et al., 2022). There were two versions to this prompt. The first simply asked the model devise a plan and solve step-by-step. The second version (PS+) extended the prompt by specifically asking the prompt to extract relevant variables and their corresponding numerals and to calculate intermediate results. We used PS+ on our experiments. One difference between PS+ and QAP is that PS+ prompt is more specific to math datasets - as it instructs to extract variables, intermediate results etc, whereas QAP is more general. Also, PS+ prompts the model to understand the problem, but it is not clear if model should output anything specific to the question itself, but QAP explicitly instructs the model to explain the problem in n words.
@@ 8 Conclusion
In this paper, we explored the approach of question analysis prompting to improve LLM accuracy across math and commonsense reasoning. The ability of this prompting method to perform well in diverse model types and tasks difficulty and type of tasks seems promising. To our best understanding, QAP is the first zero-shot prompt to introduce adaptability with a configurable parameter. We plan to extend this work further by combining QAP with other prompt strategies,  applying decoding strategies and evaluating multi-modal tasks.
@@ 10 Ethics
We experimented on three arithmetic datasets: GSM8K (Cobbe et al., 2021), AQuA (Ling et al., 2017), and AGIEval SAT Math (Zhong et al., 2023). For commonsense reasoning, used StrategyQA (Geva et al., 2021). GSM8K use the MIT License code, while AQUA and StrategyQA use the Apache-2.0 code. QAP and the prompts used in this work do not jeopardize the safety of others. They do not include any wording which may deem offensive to any individual or group.
@@ A.3 Example Explanations
Figure 3: Examples of QAP inducing explanations of the question on GSM8K, AQuA, and StrategyQA. The prompts include QAP50, QAP150, QAP50 respectively. Pink highlights key phrases (math reasoning) and orange highloghts represents useful background information (commonsense reasoning).
@@ A.5 Large value of n for simple problems hurts the performance
A.6 Word Counts for all datasets with GPT 3.5 and GPT 4
Figure 7: Median word counts in response for all datasets using GPT 3.5 Turbo and GPT 4 Turbo

### v2

@@ Abstract
Although LLMs have the potential to transform many fields, they still underperform humans in reasoning tasks. Existing methods induce the model to produce step-by-step calculations, but this research explores the question: Does making the LLM analyze the question improve its performance? We propose a novel prompting strategy called Question Analysis Prompting (QAP), in which the model is prompted to explain the question in n words before solving. The value of n influences the length of response generated by the model. QAP is evaluated on GPT-3.5 Turbo and GPT-4 Turbo on arithmetic datasets GSM8K, AQuA, and SAT and commonsense dataset StrategyQA. QAP is compared with other state-of-the-art prompts including chain-of-thought (CoT), Plan and Solve Prompting (PS+) and Take A Deep Breath (TADB). QAP outperforms all state-of-the-art prompts on AQuA and SAT datasets on both GPT-3.5 and GPT-4. QAP consistently ranks among the top-2 prompts on 75% of the tests. A key factor of QAP performance can be attributed to response length, where detailed responses are beneficial when answering harder questions, but can negatively affect easy questions.
@@ 1 Introduction
Large language models (LLMs) have recently shown rapid improvement across a host of standard natural language processing (NLP) tasks, including arithmetic, commonsense and symbolic reasoning. (Brown et al., 2020) Although these models show improved ability to understand and generate text (OpenAI, 2023), their performance can still be further improved. One solution is to encourage the model to think step-by-step. Using chain-of-thought prompting (Wei et al., 2022), LLMs are given Q&A exemplars which are designed to elicit a structured step-by-step response from the model. Many newly developed strategies meant to improve LLM performance have been focused on sophisticating the model’s step-by-step calculation (Gu et al., 2023). Despite SoTA prompts’ remarkable success across various tasks, their accuracies can still be further improved. In this work, we explore ways to improve the model reasoning not only in the answer steps, but also how the model interprets the question itself. By making the model to explicitly interpret the question, we maximize its understanding of the question and minimize missed key information. This paper introduces Question-Analysis Prompting (QAP), a simple zero-shot prompting strategy that induces the model to first explain the question before solving. We include a configurable parameter within the prompt to examine how different word counts affect the quality of a model’s response.
@@ 2 Prompt Design
The key principle behind QAP is that the model should reiterate the problem in its own words before solving. The benefit is that the model will be able to first think about what task it is trying to solve before it pursues the answer. Another principle is that we should be able to control how much the model explains so that we can adapt the prompt to different model sizes and problem complexities. The specific prompt used is as follows:
@@ 3 Prompt Impact
In Figure 1, we highlight the structure of a standard QAP output. First, the model breaks down the question in its own words and provides detailed analysis on each event. Many of the steps highlighted in the explanation were shown in the calculation section. Compared to the CoT output, QAP encourages more sophistication in its response and thus reaches the correct answer.
@@ 4.1 Benchmarks
We evaluate the effectiveness of QAP on three arithmetic reasoning datasets. These include grade-school math questions from GSM8K (Cobbe et al., 2021), algebraic word problems from AQuA (Ling et al., 2017), and SAT math problems from AGIEval (Zhong et al., 2023). For commonsense reasoning, we evaluate on open-domain questions that require implicit reasoning, from StrategyQA (Geva et al., 2021). We evaluate on the test sets of all benchmarks.
@@ 4.2 Models
We specifically choose our models to observe the prompts’ impacts across differences in model size. The smaller model is GPT-3.5 Turbo with version gpt-3.5-turbo-0613. Our larger model is GPT-4 Turbo with version gpt-4-1106-preview (OpenAI, 2023). For both of the models we used the OpenAI API  [^1] https://platform.openai.com/docs/api-reference/chat for running our experiments. The temperature and Top-K sampling was set to 0 to avoid randomness and keep consistency in the model’s responses.
@@ 4.3 Prompts
For all datasets and models, we experiment with different variations of QAP. We utilize QAP25, QAP50, QAP100, QAP150, and QAP200. We compare the performance of QAP with the baseline (no prompt). Additionally we compare QAP with two different zero-shot prompts: TADB - "Take a deep breath and work on this problem step-by-step" (Yang et al., 2023) and PS+ (Plan and Solve Plus) (Wang et al., 2023). Finally we also compare QAP with 8-shot chain-of-thought prompting.
@@ 4.4 Results
The results for GPT-3.5 Turbo and GPT-4 Turbo are shown in  Table 1 and  Table 2 respectively. General word counts are shown in Figure 7.
Arithmetic Reasoning: On GPT-3.5 Turbo, a variant of QAP is the top performer in 2 out of 3 arithmetic tasks. QAP shows significant gains on AQuA and SAT. With GPT-4 Turbo, QAP performs the best in the same 2 out of 3 arithmetic tasks. This suggests that QAP may be more beneficial on questions involving algebraic and higher-level problem solving.
@@
Table 1: Results for GPT-3.5 Turbo
@@
Table 2: Results for GPT-4 Turbo.
Commonsense Reasoning:. On StrategyQA, QAP consistently performs second-best when compared to other prompts. On both models, QAP25 is the highest QAP performer. This suggests that fewer-word explanations benefit commonsense reasoning. This is because too much explanation can cause the model to confuse a simple answer (shown in Figure 6. While there is a decline in performance as $n$ increases on the 3.5 model, the larger GPT-4 Turbo model yields similar performances across all QAP variants.
@@ 5 Analysis
Question Difficulties Based On Baseline Performance: Within a given dataset, the difficulty of the individual question may vary. We propose a method to measure question difficulty based on performance with the baseline prompt. If the model can answer the problem correctly with the baseline prompt, then we consider the question to be easy; otherwise the question is hard. We analyze the performance of different prompts across "easy" and "hard" questions. Table 3 and Table 4 show that QAP consistently outperforms other prompts in the “hard” category.
Impact Of Word Counts On Question Difficulties: QAP generates higher word counts for both “easy" and “hard" questions ( Table 5 and  Table 6 ), despite performing lower on “easy” questions. Although more step-by-step thought processes are encouraged to avoid mistakes during reasoning, this suggests that over-explanation can negatively impact the model (also shown in  Figure 5). Thus, the most suitable word count to solve a problem will vary from task to task; longer explanations are best suited to more complicated questions for which baseline prompting fails.
Downsides Of Smaller QAPs: Despite high performance on StrategyQA, QAP25 performs poorly on arithmetic datasets (mostly SAT and AQuA) using GPT-3.5 Turbo. Due to a small value of n, the model outputs are unfinished responses (i.e. the model stops midway through its reasoning steps) (shown in Figure 8). On SAT math, 51% of responses were incomplete for QAP25. On AQuA, 19% of responses were incomplete for QAP25.
Figure 2: We consider difficulty of the problem based on baseline’s results. E.g., an incorrect answer is “hard” and a correct answer is “easy”. Left chart shows accuracy within each difficulty. Right chart shows mean (average) word count for within each difficulty. All results for each prompt are shown in Table 6 and Table 4
@@ 6 Additional Studies
Placement of the prompt: In this evaluation, we studied the impact of prompt placement on performance using GSM8K dataset. Two options for prompt placement were considered: Q_Begin - adding the prompt before the question, and Q_End - adding the prompt after the question. Both placements provided similar results on GPT-3.5 and GPT-4. Results shown in the rest of the paper are based on Q_End.
No N Constraint: To test the effectiveness of adding the value of N, we first examine the prompt with just the phrase: "Explain this problem to me. Then solve for the answer". However, the model does not explain the question completely and in most cases directly starts solving the question. Its responses are no different than a response which used no given prompt. This shows that explicitly stating the minimum amount of words required is more likely to induce the model to explicitly generate an explanation of the question.
@@ 7 Related Work
Chain of Thought: Chain-of-thought reasoning is a notable few-shot (zero-shot also exists (Yang et al., 2023) example in which the model is shown how to express its reasoning steps (Wei et al., 2022). This approach was highly effective as the model would replicate these exemplars, and their accuracies improved drastically. CoT encouraged the model to think step-by-step, and this concept would be a repeating theme among other zero-shot counterparts.
@@
Plan and Solve Prompting Plus: Another zero-shot prompt is Plan-and-Solve Prompting (Wang et al., 2022). There were two versions to this prompt. The first simply asked the model devise a plan and solve step-by-step. The second version (PS+) extended the prompt by specifically asking the prompt to extract relevant variables and their corresponding numerals and to calculate intermediate results. We used PS+ on our experiments. One difference between PS+ and QAP is that PS+ prompt is more specific to math datasets since it instructs to extract variables, intermediate results, etc., whereas QAP is more general. Also, PS+ prompts the model to understand the problem, but it is not clear if model should output anything specific to the question itself. In contrast, QAP explicitly instructs the model to explain the problem in n words.
@@ 8 Conclusion
In this paper, we explored the approach of Question-Analysis Prompting to improve LLM accuracy across math and commonsense reasoning. The prompt focuses on how the model interprets the task given, and whether restating the question in its own words can further sophisticate its answer steps. The ability of this prompting method to perform well in diverse model types, tasks difficulty, and type of tasks seems promising. We plan to extend this work further by combining QAP with other prompt strategies,  applying decoding strategies and evaluating multi-modal tasks.
@@ 10 Ethics
We experiented on three arithmetic datasets: GSM8K (Cobbe et al., 2021), AQuA (Ling et al., 2017), and AGIEval SAT Math (Zhong et al., 2023). For commonsense reasoning, used StrategyQA (Geva et al., 2021). GSM8K use the MIT License code, while AQUA and StrategyQA use the Apache-2.0 code. QAP and the prompts used in this work do not jeopardize the safety of others. They do not include any wording which may deem offensive to any individual or group.
@@ A.3 Example Explanations
Figure 3: Examples of QAP inducing explanations of the question on GSM8K, AQuA, and StrategyQA. The prompts include QAP50, QAP150, QAP50 respectively. Pink highlights key phrases (math reasoning) and orange highlights represents useful background information (commonsense reasoning).
@@ A.5 Large value of n for simple problems hurts the performance
A.6 Word Counts for all datasets with GPT-3.5 and GPT-4
Figure 7: Median word counts in response for all datasets using GPT-3.5 Turbo and GPT-4 Turbo